Two trains each of length 100 m moving parallel towards each other at speed 72km/h and 36km/h respectively. In how much time will they cross each other

Grade:11

2 Answers

Latika Leekha
8 years ago
Hello student,
Both the trains are of lenght 100m.
Hence, total distance = 100 + 100 = 200 m
Relative speeds = 72 + 36 = 108 km/h = 108. 1000/ 3600 m/s
= 30 m /s
Hence, time taken = Distance/ Speed
= 200/ 30
= 6.67 sec
vansh chaudhary
6 years ago
since the question states that train are running towards eah other that means their relative speed will be diffrence b/w them.
hence, relative speed = 72-36= 36 km/hr = 36 *(5/18) = 10m/sec
now time taken= dist/speed
200/10
​=20 secs